Write the equivalent rational numbers of -5/3 with numerator is -15
 

Grade:7

1 Answers

Yogendra Kumar
2 years ago
Hi

the equivalent rational numbers of -5/3 with numerator is -15

\frac{-5}{3}=\frac{-5\times 3}{3\times3}=\frac{-15}{9}

Thanks and Regards